# The Procrustean Bed of Time Series: The Optimization Bias of Point-wise Loss

### v1

@@ 6.3 Overall Performance Analysis
Table 1 presents the comprehensive evaluation of proposed loss function deployed on PatchTST against a wide array of SOTA baselines on long-term forecasting tasks. As evidenced by the results, our Harmonized $\ell_{p}$ Norm framework consistently achieves superior performance, securing the champion (bold) or runner-up (underlined) position across the majority of datasets and prediction horizons against the SOTA Transformer-, MLP-, and CNN-based algorithms. We argue that although $\mathcal{L}_{\text{Harm},\ell_{p}}$ does not achieve the optimal performance in some dataset, but it achieves significant improvement compared with basis model PatchTST.
6.4 Ablation Studies on Optimization Objectives
To verify the effectiveness and universality of our proposed Principled Debiasing Program, we conducted ablation experiments by replacing the standard Temporal Mean Squared Error ($\mathcal{L}_{\text{TMSE}}$) with our proposed Harmonized $\ell_{1}$ Norm ($\mathcal{L}_{\text{Harm},\ell_{1}}$), while the former serves as the default training objective for the vast majority of models. As evidenced by quantitative results are summarized in Table 2, models trained with our $\mathcal{L}_{\text{Harm},\ell_{1}}$ consistently outperform those trained with $\mathcal{L}_{\text{TMSE}}$ across almost all datasets, horizons, and architectures, showing a consist performance improvement regardless of the underlying model structure.
@@ 7 Conclusion
This work shifts the focus from network expressiveness to the fundamental training paradigm. By scrutinizing the point-wise i.i.d. assumption through a rigorous information-theoretic lens, we formalized the Optimization Bias, unveiled the Paradigm Paradox, and derived the unified closed-form quantification of the non-deterministic EOB. These insights motivated a Principled Debiasing Program, which we empirically validated to surpass standard baselines using temporal MSE. Ultimately, we hope to dispel the haze of AI alchemy and and nurture its sustainable growth from the ground up.
@@
Table of Appendix Contents
@@
The proof of bounds of expectation of optimization bias is divided into two parts, corresponding to the left and right inequalities. We use the definition of entropy $H(a_{t})$, mutual information $I(a_{t};a_{1:t-1})$ and $I(a_{1:T})$:
@@
The proof is based on the following basic assumptions:
  • (1) Determination: the deterministic component $v_{t}$ is perfectly predictable from its past, i.e., $v_{t}=f(v_{1:t-1})$, such that $p(v_{t}|v_{1:t-1})$ is a Dirac delta function $\delta(x)$, and $H(v_{t}|v_{1:t-1})=$-$\infty$.
  • (2) Independence: The processes $\{v_{t}\}$ and $\{z_{t}\}$ is independent.
  • (3) Information Equivalence: The information contained in the history $x_{1:t-1}$ is equivalent to the information contained in $\{v_{1:t-1},z_{1:t-1}\}$.
Then the theorem is equivalent to the inequality about mutual information $I(z_{t})$, $I(x_{t})$, and $I(v_{t})$:
$I(z_{1:T})\leq I(x_{1:T})\leq I(v_{1:T}).$ (32)
A.1 Part 1: Proof of $I(z_{1:T})\leq I(x_{1:T})$
To prove this, we show that $I(z_{t};z_{1:t-1})\leq I(x_{t};x_{1:t-1})$ for any $t\in[2,T]$.
First, we establish that the conditional entropy of $x_{t}$ given its past is equal to the conditional entropy of $z_{t}$ given its past.
@@ A.1 Part 1: Proof of $I(z_{1:T})\leq I(x_{1:T})$
$=H(v_{t}+z_{t}|v_{1:t-1},z_{1:t-1})$ (By Condition 3: Information Equivalence)
$=H(z_{t}|v_{1:t-1},z_{1:t-1})$ (By Condition 1: Determination)
$=H(z_{t}|z_{1:t-1})$ (By Condition 2: Independence)
Next, we compare the marginal entropy of $x_{t}$ and $z_{t}$.
$H(x_{t})$ $=H(v_{t}+z_{t})$ (By Decomposition) (34)
$\geq H(z_{t})$ (By Condition 2: Independence)
We now compare the mutual information terms:
@@
Thus, $I(x_{t};x_{1:t-1})\geq I(z_{t};z_{1:t-1})$. Summing over $t=2$ to $T$ yields $I(x_{1:T})\geq I(z_{1:T})$.
A.2 Part 2: Proof of $I(x_{1:T})\leq I(v_{1:T})$
We analyze the mutual information of the deterministic component $\{v_{t}\}$:
@@ A.2 Part 2: Proof of $I(x_{1:T})\leq I(v_{1:T})$
By Condition 1 (Determination), the process is perfectly predictable, so its conditional entropy is $H(v_{t}|v_{1:t-1})=$-$\infty$. Meanwhile, assuming $\{v_{t}\}$ is not a fixed constant, its marginal entropy $H(v_{t})$ is a finite value. Therefore, the mutual information for a single step is:
$I(v_{t};v_{1:t-1})=H(v_{t})-(-\infty)=\infty.$ (37)
The total correlation is a sum of these terms, which is also infinite:
$I(v_{1:T})=\sum_{t=2}^{T}I(v_{t};v_{1:t-1})=\infty.$ (38)
We analyze the mutual information of the mixed process $\{x_{t}\}$:
$I(x_{t};x_{1:t-1})=H(x_{t})-H(x_{t}|x_{1:t-1}).$ (39)
From Eq. (33), we proved $H(x_{t}|x_{1:t-1})=H(z_{t}|z_{1:t-1})$. Because $\{z_{t}\}$ is a purely non-deterministic process, it has an innovation component with non-zero variance. This means its conditional entropy $H(z_{t}|z_{1:t-1})$ is a finite value (and strictly $\infty$). The marginal entropy $H(x_{t})$ is also assumed to be finite for a non-pathological process. Therefore, $I(x_{t};x_{1:t-1})$ is the difference between two finite numbers, and is itself finite. The total correlation $I(x_{1:T})$ is a sum of finite terms, which is also finite.
Since $I(x_{1:T})$ is finite and $I(v_{1:T})$ is infinite, the inequality $I(x_{1:T})\leq I(v_{1:T})$ trivially holds.
@@
Notion: The theoretical infinite upper bound of EOB derived from pure deterministic components exists solely as an abstraction. In practice, this bound is constrained by inevitable background noise during data sampling and the finite precision of numerical computing.
@@
For $AR(p)$ process, we can specify the exact conditional and marginal distributions required to calculate $\mathcal{B}_{z}$:
  • Conditional Distribution: The distribution of $x_{t}$ given its past is determined directly by the model’s structure. Due to the Markov property of $AR(p)$ model, this simplifies to conditional only on the $p$ preceding time steps.
    $p(z_{t}|z_{1:t-1})\sim\mathcal{N}(\mu_{t},\sigma_{t}^{2})$ (40)
    where $\mu_{t}=c+\sum_{i=1}^{p}\phi_{i}z_{t-i}$ and $\sigma_{t}^{2}=\sigma_{\epsilon}^{2}$.
  • Marginal Distribution: For a stationary $AR(p)$ process, the unconditional distribution of any observation $z_{t}$ is also Gaussian, with a constant mean and variance.
@@
The summation starts from $t=p+1$ as this is the first time point for which the conditional distribution $p(z_{t}|z_{1:t-1})$ is fully defined by the $p$ preceding observations, simplifying the analysis by focusing on the steady-state behavior of the process.
Substituting these Gaussian PDFs into the definition of $\mathcal{B}_{z}$ (Eq. (3)), we obtain:
$\mathcal{B}_{z}$ $=\sum_{t=p+1}^{T}\log\frac{p(z_{t}|z_{1:t-1})}{p(z_{t})}$ (43)
$=\sum_{t=p+1}^{T}\left[\frac{1}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{t}^{2}}-\frac{(z_{t}-\mu_{t})^{2}}{2\sigma_{t}^{2}}+\frac{(z_{t}-\mu)^{2}}{2\sigma_{z}^{2}}\right]$
$=\sum_{t=p+1}^{T}\left[\frac{1}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}-\frac{\epsilon_{t}^{2}}{2\sigma_{\epsilon}^{2}}+\frac{(z_{t}-\mu)^{2}}{2\sigma_{z}^{2}}\right].$
The expected optimization bias, $\mathbb{E}[\mathcal{B}_{z}]$, reveals the average magnitude of this systemic error. The expectation of the fluctuation term cancel out, since $\mathbb{E}[(z_{t}-\mu)^{2}]=\operatorname{Var}(z_{t})=\sigma_{z}^{2}$ and $\mathbb{E}[\epsilon_{t}^{2}]=\sigma_{\epsilon}^{2}$. This leaves a remarkably clean result:
$\mathbb{E}[\mathcal{B}]=\frac{T-p}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}=\frac{T-p}{2}\log\frac{1}{1-\sum_{i=1}^{p}\phi_{i}\rho_{i}}.$ (44)
@@ Discussion.
The proof begins with Szegő’s First Limit Theorems, a powerful result concerning the determinants of Toeplitz matrix. For a covariance-stationary process, the correlation matrix $R$ is a Toeplitz matrix, with elements $R_{ij}=\rho_{|i-j|}$. The theorem connects the asymptotic behavior of the determinant $|R|$ to the power spectral density, $f(\omega)$, of the process:
$\lim_{T\to\infty}\frac{1}{T}\log|R|=\frac{1}{2\pi}\int_{-\pi}^{\pi}\log f(\omega)d\omega.$ (62)
For an $AR(p)$ process, the power spectral density is given by:
$f(\omega)=\frac{\sigma_{\epsilon}^{2}}{2\pi\left|1-\sum_{k=1}^{p}\phi_{k}e^{-ik\omega}\right|^{2}}.$ (63)
Applying Kolmogorov’s Formula, the integral of the log-spectrum for any stationary, non-deterministic process simplifies to the log-ratio of the innovation variance ($\sigma_{\epsilon}^{2}$) to the marginal variance ($\sigma_{z}^{2}$):
$\frac{1}{2\pi}\int_{-\pi}^{\pi}\log f(\omega)d\omega=\log\sigma_{\epsilon}^{2}-\log\sigma_{z}^{2}=\log\left(\frac{\sigma_{\epsilon}^{2}}{\sigma_{z}^{2}}\right).$ (64)
Combining these results, we find the limit:
$\lim_{T\to\infty}\log|R|^{1/T}=\log\left(\frac{\sigma_{\epsilon}^{2}}{\sigma_{z}^{2}}\right).$ (65)
Exponentiating both sides yields the final result:
@@ Appendix E Non-linear Modeling of Non-deterministic EOB
$\mathbb{E}[\mathcal{B}_{z}]$ $=\mathbb{E}\left[\log\frac{p(z_{1:T})}{\prod_{t=1}^{T}p(z_{t})}\right]$ (79)
$=\mathbb{E}[\log p(z_{1:T})]-\sum_{t=1}^{T}\mathbb{E}[\log p(z_{t})]$

### v2

@@ 6.4 Overall Performance Analysis
Table 1 and Table 2 presents a part of comprehensive evaluation of our proposed $\mathcal{L}_{\text{Harm},\ell_{p}}$ deployed on iTransformer backbone, compared against SOTA Transformer-, MLP-, and CNN-based baselines.
Superiority in Long-term Forecasting. Detailed results are provided in Appendix J.3 conducted on 11 benchmarks against 13 SOTA baselines. As shown in Table 1, our novel loss function demonstrates overwhelming superiority, particularly in the MAE metric. Out of 44 evaluated cases (11 datasets $\times$ 4 horizons), $\mathcal{L}_{\text{Harm},\ell_{p}}$ secures 17 champions in MSE and 31 in MAE. Notably, our loss achieves 90.9% Top-2 coverage (40 out of 44 cases) in MAE, significantly outperforming the second-best model, FreDF. This consistent lead across diverse datasets underscores the exceptional robustness of our harmonized optimization objective.
Excellence in Missing Data Imputation. Detailed results are provided in Appendix J.4 conducted on 9 benchmarks against 9 baselines. In Table 2, our method maintains its competitive edge by securing 13 champions in both MSE and MAE out of 36 cases. More impressively, we achieve a 91.7% Top-2 coverage (33 out of 36 cases) in MSE, effectively dominating the imputation benchmarks. The framework’s ability to precisely recover missing values across various missing rates ($p_{\text{miss}}$) further validates its capability to reconstruct the underlying data manifold.
Optimization-Driven Improvements. Crucially, since our loss function is built upon the iTransformer backbone, the direct comparison with the vanilla iTransformer reveals the intrinsic value of our contribution. Our loss yields consistent and significant error reductions across nearly all scenarios. This empirical evidence indicates that the performance leap stems not merely from the architectural design, but from the superior optimization objective via $\mathcal{L}_{\text{Harm},\ell_{p}}$.
Plug-and-play Versatility. To verify the generalization of our debiasing strategy, we extend $\mathcal{L}_{\text{Harm},\ell_{p}}$ to various models. As detailed in Appendix J.5, the uniform gains achieved across various architectures underscore the model-agnostic nature of EOB theory.
6.5 Analysis of Structural Orthogonality Methods
To quantitatively evaluate and guide the selection of optimal unitary transformations across diverse datasets, we introduce a suite of four diagnostic metrics: Off-Diagonal Energy (ODE) ratio, Spearman Mean, Eigen-Entropy, and Distance to Identity. These metrics collectively characterize the degree of structural orthogonality and information redistribution achieved by different unitary bases. Specifically, they measure the extent to which a transformation can convert complex temporal dependencies into a structurally orthogonal representation. (See Appendix J.6)
@@ 7 Conclusion
This work shifts the focus from network expressiveness to the fundamental training paradigm. By scrutinizing the point-wise i.i.d. assumption through a rigorous information-theoretic lens, we formalize the Optimization Bias, unveil the Paradigm Paradox, and derive a unified closed-form quantification of the non-deterministic EOB as the lower bound of EOB. These insights motivated a Principled Debiasing Program, which we empirically validated to significantly surpass standard baselines using temporal MSE on forecasting and imputation tasks. Furthermore, the proposed loss function demonstrates consistent improvements across diverse models and benchmarks. Ultimately, we hope to dispel the haze of AI alchemy and nurture its sustainable growth from the ground up.
@@
Detailed Appendix Directory
@@
The proof establishes the bounds for the Expectation of Optimization Bias. To accommodate the potentially singular nature of deterministic components (e.g., Dirac delta functions), we explicitly frame this analysis within the domain of continuous random variables. Accordingly, all entropy terms $H(\cdot)$ refer to differential entropy.
We utilize the definitions of differential entropy $H(a_{t})$, mutual information $I(a_{t};a_{1:t-1})$, and total correlation $I(a_{1:T})$:
@@
The proof relies on the following formalized assumptions regarding the decomposition $x_{t}=v_{t}+z_{t}$:
Assumption A.1.
(Continuous Determinism) The deterministic component $v_{t}$ is a continuous variable governed by a deterministic function of its past, i.e., $v_{t}=f(v_{1:t-1})$. Its conditional distribution approaches a Dirac delta function, $p(v_{t}|v_{1:t-1})\to\delta(v_{t}-f(v_{1:t-1}))$, implying its conditional differential entropy diverges: $H(v_{t}|v_{1:t-1})\to$-$\infty$.
Assumption A.2.
(Independence) The deterministic process $\{v_{t}\}$ and the stochastic process $\{z_{t}\}$ are statistically independent.
Assumption A.3.
(Structural Identifiability) We assume the mixture is structurally identifiable given the infinite history. Specifically, the information contained in the observed history $x_{1:t-1}$ is sufficient to recover the underlying components $\{v_{1:t-1},z_{1:t-1}\}$. This implies the sigma-algebra generated by $x_{1:t-1}$ is equivalent to that of the joint pair, justifying $H(\cdot|x_{1:t-1})=H(\cdot|v_{1:t-1},z_{1:t-1})$.
A.1 Part 1: Proof of Lower Bound $I(z_{1:T})\leq I(x_{1:T})$
To prove the lower bound, it suffices to show that $I(z_{t};z_{1:t-1})\leq I(x_{t};x_{1:t-1})$ for any $t\in[2,T]$.
Step 1: Equivalence of Conditional Entropy. We first establish that the uncertainty of the current observation $x_{t}$ given its history is entirely due to the stochastic component $z_{t}$.
@@ A.1 Part 1: Proof of Lower Bound $I(z_{1:T})\leq I(x_{1:T})$
$=H(v_{t}+z_{t}|v_{1:t-1},z_{1:t-1})$ (By Assumption A.3: Identifiability)
[…]
$=H(z_{t}|z_{1:t-1})$ (By Assumption A.2: Independence)
Notably, step 3 utilizes the translation invariance of differential entropy, $H(z_{t}+c|\cdot)=H(z_{t}|\cdot)$.
Step 2: Comparison of Marginal Entropy. Next, we compare the marginal entropies. For independent continuous variables, the entropy of a sum is at least the entropy of the individual components (a consequence of the entropy power inequality context):
$H(x_{t})$ $=H(v_{t}+z_{t})$ (31)
$\geq H(z_{t})$ (By Assumption A.2: Independence)
Intuitively, adding the independent variation of $v_{t}$ to $z_{t}$ cannot decrease the uncertainty (entropy) of the resulting distribution.
Step 3: Combining Results. Comparing the mutual information terms:
@@
Summing over $t$ yields $I(x_{1:T})\geq I(z_{1:T})$.
A.2 Part 2: Proof of Upper Bound $I(x_{1:T})\leq I(v_{1:T})$
We analyze the limiting behavior of the deterministic component’s mutual information.
@@ A.2 Part 2: Proof of Upper Bound $I(x_{1:T})\leq I(v_{1:T})$
By Assumption A.1, as the conditional distribution approaches a Dirac delta, the conditional differential entropy $H(v_{t}|v_{1:t-1})\to$-$\infty$. Assuming the marginal distribution of the deterministic trend $\{v_{t}\}$ spans a non-singular range (i.e., it is not a constant value for all time), its marginal entropy $H(v_{t})$ is finite.
Consequently, the mutual information for the deterministic component diverges:
$I(v_{t};v_{1:t-1})\to\infty\implies I(v_{1:T})\to\infty.$ (34)
In contrast, consider the mixed process $\{x_{t}\}$. From Eq. (30), we have $H(x_{t}|x_{1:t-1})=H(z_{t}|z_{1:t-1})$. Since $\{z_{t}\}$ contains a non-degenerate stochastic innovation ($\sigma_{\epsilon}^{2}>0$), its conditional differential entropy is finite (specifically, related to $\log\sigma_{\epsilon}$). Similarly, the marginal entropy $H(x_{t})$ is finite. Thus, $I(x_{1:T})$ represents a sum of finite terms and is bounded.
Since $I(x_{1:T})<\infty$ and $I(v_{1:T})\to\infty$, the inequality $I(x_{1:T})\leq I(v_{1:T})$ holds in the limit of perfect determinism.
@@
To strictly quantify the EOB for an $AR(p)$ process, we must account for the dependence structure. The exact conditional distribution $p(z_{t}|z_{1:t-1})$ varies depending on whether $t\leq p$ (transient phase) or $t>p$ (steady-state phase).
We decompose the total EOB into two components: the initial transient terms and the steady-state accumulation:
$\mathcal{B}_{z}=\sum_{t=2}^{T}\log\frac{p(z_{t}|z_{1:t-1})}{p(z_{t})}=\underbrace{\sum_{t=2}^{p}\log\frac{p(z_{t}|z_{1:t-1})}{p(z_{t})}}_{\text{Transient Term }(\mathcal{B}_{\text{tran}})}+\underbrace{\sum_{t=p+1}^{T}\log\frac{p(z_{t}|z_{1:t-1})}{p(z_{t})}}_{\text{Steady-State Term }(\mathcal{B}_{\text{ss}})}$ (35)
1. Steady-State Phase ($t>p$):
For $t>p$, the conditional distribution is fully determined by the $p$ preceding observation. Due to the Markov property and stationarity:
  • Conditional Distribution:
    $p(z_{t}|z_{1:t-1})\sim\mathcal{N}(\mu_{t},\sigma_{\epsilon}^{2})$ (36)
    where $\mu_{t}=c+\sum_{i=1}^{p}\phi_{i}z_{t-i}$. Note that the conditional variance stabilizes to the innovation variance $\sigma_{\epsilon}^{2}$.
  • Marginal Distribution:
@@ 1. Steady-State Phase ($t>p$):
Substituting these into the steady-state sum $\mathcal{B}_{\text{ss}}$:
$\mathcal{B}_{\text{ss}}$ $=\sum_{t=p+1}^{T}\left[\frac{1}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}-\frac{(z_{t}-\mu_{t})^{2}}{2\sigma_{\epsilon}^{2}}+\frac{(z_{t}-\mu)^{2}}{2\sigma_{z}^{2}}\right].$ (39)
Taking the expectation $\mathbb{E}[\mathcal{B}_{\text{ss}}]$, the fluctuation terms cancel out (since $\mathbb{E}[(z_{t}-\mu_{t})^{2}]=\sigma_{\epsilon}^{2}$ and $\mathbb{E}[(z_{t}-\mu)^{2}]=\sigma_{z}^{2}$), yielding:
$\mathbb{E}[\mathcal{B}_{\text{ss}}]=\frac{T-p}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}$ (40)
2. Transient Phase ($t<p$):
For the initial time steps $t<p$, the conditional variance $\sigma_{t}^{2}=\operatorname{Var}(z_{t}|z_{1:t-1})$ is not equal to $\sigma_{\epsilon}^{2}$, but rather lies in the interval $\sigma_{\epsilon}^{2},\sigma_{z}^{2}]$ as the process uncertainty reduces from the marginal to conditional level. The sum of these initial terms constitutes a constant $C_{p}$ that depends only on the model parameters ($p,\phi$) but is independent of $T$.
$\mathbb{E}[\mathcal{B}_{\text{tran}}]=\sum_{t=2}^{p}\frac{1}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{t}^{2}}=C_{p}$ (41)
Total Expected EOB:
Combining both parts, the total expected EOB is:
$\mathbb{E}[\mathcal{B}_{z}]=\frac{T-p}{2}\log\left(\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}\right)+C_{p}=\frac{T-p}{2}\log\left(\frac{1}{1-\sum_{i=1}^{p}\phi_{i}\rho_{i}}\right)+C_{p}.$ (42)
Since $C_{p}$ is constant with respect to $T$, for large sequence lengths ($T\gg p$), the EOB is dominated by the steady-state linear growth. Thus, we focus on the asymptotic rate:
$\mathbb{E}[\mathcal{B}_{z}]\approx\frac{T}{2}\log\frac{\sigma_{z}^{2}}{\sigma_{\epsilon}^{2}}\quad\text{as }T\to\infty.$ (43)
@@ Discussion.
The proof utilizes Szegő’s First Limit Theorem, which relates the asymptotic determinant of a Toeplitz matrix to the geometric mean of its generating function. For the correlation matrix $R$, the generating function is the normalized power spectral density, $g(\omega)=f(\omega)/\sigma_{z}^{2}$, where $f(\omega)$ is the PSD of the process.
The theorem states:
$\lim_{T\to\infty}\frac{1}{T}\log|R|=\frac{1}{2\pi}\int_{-\pi}^{\pi}\log\left(\frac{f(\omega)}{\sigma_{z}^{2}}\right)d\omega.$ (61)
By Kolmogorov’s Formula, the geometric mean of the process PSD $f(\omega)$ is equal to the innovation variance $\sigma_{\epsilon}^{2}$:
$\exp\left(\frac{1}{2\pi}\int_{-\pi}^{\pi}\log f(\omega)d\omega\right)=\sigma_{\epsilon}^{2}.$ (62)
Substituting this into the limit equation:
$\lim_{T\to\infty}|R|^{1/T}$ $=\exp\left(\frac{1}{2\pi}\int_{-\pi}^{\pi}\left(\log f(\omega)-\log\sigma_{z}^{2}\right)d\omega\right)$ (63)
$=\frac{\exp\left(\frac{1}{2\pi}\int_{-\pi}^{\pi}\log f(\omega)d\omega\right)}{\sigma_{z}^{2}}$
$=\frac{\sigma_{\epsilon}^{2}}{\sigma_{z}^{2}}.$
This confirms the relationship with the Signal-to-Noise Ratio (SSNR):
$\lim_{T\to\infty}|R|^{1/T}=\frac{1}{\textit{SSNR}}.$ (64)
@@ Appendix E Non-linear Modeling of Non-deterministic EOB
$\mathbb{E}[\mathcal{B}_{z}]$ $=\mathbb{E}\left[\log\frac{p(z_{1:T})}{\prod_{t=1}^{T}p(z_{t})}\right]=\mathbb{E}[\log p(z_{1:T})]-\sum_{t=1}^{T}\mathbb{E}[\log p(z_{t})]$ (77)